Question 12 4/4 pts A solution has a pH value of 3.84. What is the pOH for this solution? CALCULATIONS MUST BE PROVIDED FOR F show work
Answer #1

Q12. Given : pH = 3.84

Autoionization constant of water, Kw = 1.0 x 10-14

pKw = -log(Kw)

pKw = -log(1.0 x 10-14)

pKw = 14

pOH = pKw - pH

pOH = 14 - 3.84

pOH = 10.16